# Le Wang1, Jun Wang2, Feng Deng2, Chen Zhang2, Di Zhang2, Kun Gai2

### v2

@@ Abstract
We present AudioGen-Omni — a unified approach based on multimodal diffusion transformers (MMDit), capable of generating high-fidelity audio, speech, and songs coherently synchronized with the input video. AudioGen-Omni introduces a novel joint training paradigm that seamlessly integrates large-scale video-text-audio corpora, enabling a model capable of generating semantically rich, acoustically diverse audio conditioned on multimodal inputs and adaptable to a wide range of audio generation tasks. AudioGen-Omni employs a unified lyrics-transcription encoder that encodes graphemes and phonemes from both sung and spoken inputs into dense frame-level representations. Dense frame-level representations are fused using an AdaLN-based joint attention mechanism enhanced with phase-aligned anisotropic positional infusion (PAAPI), wherein RoPE is selectively applied to temporally structured modalities to ensure precise and robust cross-modal alignment. By unfreezing all modalities and masking missing inputs, AudioGen-Omni mitigates the semantic constraints of text-frozen paradigms, enabling effective cross-modal conditioning. This joint training approach enhances audio quality, semantic alignment, and lip-sync accuracy, while also achieving state-of-the-art results on Text-to-Audio/Speech/Song tasks. With an inference time of 1.91 seconds for 8 seconds of audio, it offers substantial improvements in both efficiency and generality.
@@ 1 Introduction
Recent works on video-to-audio and video-to-speech generation have made important strides in modeling cross-modal relationships. For example, Tian et al. Tian et al. (2025b) propose a dedicated architecture for any-to-audio synthesis that supports a broad range of input modalities—including text, video, image, music, and audio. However, the model does not support speech or singing voice synthesis tasks and exhibits suboptimal alignment between audio effects and rhythmic timing. MMAudio Cheng et al. (2025) generates synchronized audio from video and/or text via multimodal joint training and a synchronization module for temporal alignment. However, it cannot synthesize speech or singing voice, limiting its scope. DualDub Tian et al. (2025a) proposes a unified framework that jointly generates background audio and speech using a multimodal encoder and cross-modal aligner for improved synchronization. However, it lacks explicit lip-sync alignment between speech and video and does not support singing voice generation. Face2Voice Kim et al. (2025) achieves superior speech synthesis by bridging the modality gap via hierarchical representation learning but lacks support for sound effects or songs, and its language naturalness requires improvement. VidMuse Tian et al. (2025c) advances video-to-music generation with a large dataset and effective alignment techniques but does not support speech or singing voice synthesis.
@@ 2 Related Work
Song Generation. Early works. Jukebox Dhariwal et al. (2020) pioneered full-song synthesis by cascading multi-scale VQ-VAEs and transformers, yet its control is limited to genre/artist tags and inference is slow. Multi-stage pipelines. Melodist Hong et al. (2003) and MelodyLM Li et al. (2023) decompose the task into text-to-MIDI, singing-voice synthesis, and vocal-to-accompaniment alignment. While they improve vocal quality, the multi-stage design complicates training/inference and their datasets are restricted to Mandarin pop. Joint modeling. SongCreator Lei et al. (2020) employs a dual-sequence LM to jointly generate vocals and accompaniment, but lacks textual control and yields muffled vocals. Freestyle Ning et al. (2006) focuses on rap generation given lyrics and beats, sacrificing melodic diversity. Yue Yuan et al. (2017) scales data and parameters with a two-stage, track-decoupled LM, achieving strong results at high cost. Commercial systems. Suno, Udio, and SeedMusic Bai et al. (2024) deliver high-fidelity songs, yet remain closed-source and provide limited controllability.
@@ 5 Conclusion
We propose AudioGen-Omni, a unified multimodal diffusion transformer that generates high-fidelity audio, speech, and songs synchronized with input video. Leveraging large-scale video-text-audio training data, it employs a unified lyrics-transcription encoder and a novel joint attention with phase-aligned positional infusion to ensure precise cross-modal alignment. By unfreezing all modalities and masking missing inputs, AudioGen-Omni overcomes limitations of text-frozen models, enabling flexible conditioning and strong generalization. It achieves state-of-the-art results on multiple audio generation tasks with efficient inference, laying the groundwork for future extensions including video generation.

### v4

# Le Wang1, Jun Wang2, Chunyu Qiang2, Feng Deng2, Chen Zhang2, Di Zhang2, Kun Gai2

@@ Abstract
We present AudioGen-Omni — a unified approach based on multimodal diffusion transformers (MMDit), capable of generating high-fidelity audio, speech, and song coherently synchronized with the input video. AudioGen-Omni introduces a novel joint training paradigm that seamlessly integrates large-scale video-text-audio corpora, enabling a model capable of generating semantically rich, acoustically diverse audio conditioned on multimodal inputs and adaptable to a wide range of audio generation tasks. AudioGen-Omni employs a unified lyrics-transcription encoder that encodes graphemes and phonemes from both song and spoken inputs into dense frame-level representations. Dense frame-level representations are fused using an AdaLN-based joint attention mechanism enhanced with phase-aligned anisotropic positional infusion (PAAPI), wherein RoPE is selectively applied to temporally structured modalities to ensure precise and robust cross-modal alignment. By unfreezing all modalities and masking missing inputs, AudioGen-Omni mitigates the semantic constraints of text-frozen paradigms, enabling effective cross-modal conditioning. This joint training approach enhances audio quality, semantic alignment, and lip-sync accuracy, while also achieving state-of-the-art results on Text-to-Audio/Speech/Song tasks. With an inference time of 1.91 seconds for 8 seconds of audio, it offers substantial improvements in both efficiency and generality.
@@ 1 Introduction
Recent works on video-to-audio and video-to-speech generation have made important strides in modeling cross-modal relationships. For example, Tian et al. Tian et al. (2025b) propose a dedicated architecture for any-to-audio synthesis that supports a broad range of input modalities—including text, video, image, music, and audio. However, the model does not support speech or singing voice synthesis tasks and exhibits suboptimal alignment between audio effects and rhythmic timing. MMAudio Cheng et al. (2025) generates synchronized audio from video and/or text via multimodal joint training and a synchronization module for temporal alignment. However, it cannot synthesize speech or singing voice, limiting its scope. DualDub Tian et al. (2025a) proposes a unified framework that jointly generates background audio and speech using a multimodal encoder and cross-modal aligner for improved synchronization. However, it lacks explicit lip-sync alignment between speech and video and does not support singing voice generation. Face2Voice Kim et al. (2025) achieves superior speech synthesis by bridging the modality gap via hierarchical representation learning but lacks support for sound effects or song, and its language naturalness requires improvement. VidMuse Tian et al. (2025c) advances video-to-music generation with a large dataset and effective alignment techniques but does not support speech or singing voice synthesis.
@@ 2 Related Work
Song Generation. Early works. Jukebox Dhariwal et al. (2020) pioneered full-song synthesis by cascading multi-scale VQ-VAEs and transformers, yet its control is limited to genre/artist tags and inference is slow. Multi-stage pipelines. Melodist Hong et al. (2003) and MelodyLM Li et al. (2023) decompose the task into text-to-MIDI, singing-voice synthesis, and vocal-to-accompaniment alignment. While they improve vocal quality, the multi-stage design complicates training/inference and their datasets are restricted to Mandarin pop. Joint modeling. SongCreator Lei et al. (2020) employs a dual-sequence LM to jointly generate vocals and accompaniment, but lacks textual control and yields muffled vocals. Freestyle Ning et al. (2006) focuses on rap generation given lyrics and beats, sacrificing melodic diversity. Yue Yuan et al. (2017) scales data and parameters with a two-stage, track-decoupled LM, achieving strong results at high cost. Commercial systems. Suno, Udio, and SeedMusic Bai et al. (2024) deliver high-fidelity song, yet remain closed-source and provide limited controllability.
@@ 5 Conclusion
We propose AudioGen-Omni, a unified multimodal diffusion transformer that generates high-fidelity audio, speech, and song synchronized with input video. Leveraging large-scale video-text-audio training data, it employs a unified lyrics-transcription encoder and a novel joint attention with phase-aligned positional infusion to ensure precise cross-modal alignment. By unfreezing all modalities and masking missing inputs, AudioGen-Omni overcomes limitations of text-frozen models, enabling flexible conditioning and strong generalization. It achieves state-of-the-art results on multiple audio generation tasks with efficient inference, laying the groundwork for future extensions including video generation.